OpenStudy (anonymous):

I am trying to solve for t in this: e^sqrt T=x^2

OpenStudy (anonymous):

I raised both sides to ln and got sqrt T =x^2

OpenStudy (anonymous):

Then I squared both sides and got T=x^2^2

OpenStudy (anonymous):

how do i go from there to T=4(lnx)^2

OpenStudy (anonymous):

or am i even doing it right??

OpenStudy (anonymous):

you are doing it right but you forgot to keep your ln (x^2) in the first step. so you already have the right form. what you should get is this: sqrt(T)=2ln(x) -- by properties of logs T=4(lnx)^2

OpenStudy (lalaly):

Sort T = lnx^2 t= (2lnx)^2 T= 4(lnx)^2

OpenStudy (anonymous):

OH i got it